OpenStudy (anonymous):

Identify the slope and y-intercept of y = –4/3x + 3 3; -4/3 3; 4/3 -4/3; 3 4/3; 3

OpenStudy (anonymous):

3; -4/3

OpenStudy (anonymous):

y=mx+c

sam (.sam.):

y=mx+c m=slope

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

m slope

OpenStudy (anonymous):

-4/3; 3 was the answer

OpenStudy (anonymous):

yep